Ok my question is about graphing of linear inequalities but I don't know what to do you see I have a question that's ask me to describe and correct the error in the graph of the inequality that o was given the question is Y is less then or equal to two X plus two then it is graphed and is shaded on the top of the line I don't see what's wrong it is doted on (-2,-2) (0,-1) (2,0) (4,1) could you please help me I'm completely lost thanks a lot :)
Found 2 solutions by stanbon, DrBeeee:
Answer by stanbon(75887)

describe and correct the error in the graph of the inequality that o was given the question is Y is less then or equal to two X plus two then it is graphed and is shaded on the top of the line I don't see what's wrong it is doted on
(-2,-2) (0,-1) (2,0) (4,1)
-----
y <= 2x+2
----
Put each of the ordered pairs into the
inequality to see if you get a true statement:
Examples:
Using (-2,2)::: 2 <= 2*(-2)+2 ; 2 <= -4+2 ; 2 <= -2;;That is false
----
Using (0,-1)::: -1 <=2(0)+2 ; -1 <=0+2 ; -1<= 2 ;; That is true.
-----
Do the same for (2,0) and for (4,1)
=====
Cheers,
Stan H.

Answer by DrBeeee(684)

First off, the four points are on the line
(1) y = x/2 - 1
And the inequality you give is
(2) y <= 2x + 2.
So I'm confused about the question.
Is (1) also an inequality? Such as
(3) y > x/2 - 1, then the solution to this pair of inequalities is a shaded area above a dotted line for (3) (the 4 points) and below (2) which is a solid line.